Corrin is 23 years older than Beth. In 9 years, Corrin will be twice as old as Beth. How old are they now?
Answer by jorel1380(3719)

C=B+23
C+9=2(B+9)
B+23+9=2B+18
14=B
37=C
Corrin is 37. Beth is 14..